# There Was Never a Bottleneck in Concept Bottleneck Models

### v1

@@ D.2 Datasets
This is a synthetic dataset with controlled variation across seven generative factors: object shape, object color, object size, camera height, background color, horizontal axis, and vertical axis. In our setup, $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors. To ensure consistency in the mapping between concepts and the target, we filter the dataset such that any combination of elements in $\left\{\bm{c},\bm{n}_{y}\right\}$ corresponds to a unique value of $\bm{y}$. All invalid combinations are removed accordingly.
@@
CelebA
: The CelebA dataset consists of over 200,000 celebrity face images, each annotated with 40 binary facial attributes. In our setup, the task variable $\bm{y}$ corresponds to binary gender classification. For each run, eight attributes are randomly selected as the concept set $\bm{c}$, while the remaining attributes are treated as nuisances $\bm{n}_{y}$. Since most attributes exhibit some degree of correlation with gender, we set $\bm{n}_{\bar{y}}$ to the empty set.
@@
: The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes. Following the approach of [9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes. These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names. In our setup, the task variable $\bm{y}$ is to classify the bird species. The concept set $\bm{c}$ consists of the attributes belonging to 7 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups. Since most attributes exhibit some correlation with the classification task, we set $\bm{n}{\bar{y}}$ to the empty set.

### v2

@@ E.2 Datasets
This is a synthetic dataset with controlled variation across seven generative factors: object shape, object color, object size, camera height, background color, horizontal axis, and vertical axis. In our setup, $\bm{y}$ corresponds to the object shape, $\bm{n}_{y}$ to the horizontal axis, $\bm{n}_{\bar{y}}$ to the vertical axis, and $\bm{c}$ to the remaining generative factors. To ensure consistency in the mapping between concepts and task nuisances and the target, we filter the dataset such that any combination of elements in $\left\{\bm{c},\bm{n}_{y}\right\}$ corresponds to a unique value of $\bm{y}$. All invalid combinations are removed accordingly.
@@
CIFAR-10
CIFAR-10 is a widely used image classification benchmark consisting of 60,000 natural images of size 32 $\times$ 32, divided into 10 classes (e.g., airplanes, automobiles, birds, cats, etc.). The dataset is split into 50,000 training and 10,000 test images, with balanced class distributions. To reduce the need for manual concept annotations, the concepts are synthetically derived following the methodology of [23]. A total of 143 attributes are extracted using GPT-3 [33]; 64 form the concept set $\bm{c}$, while the rest define the nuisance set $\bm{n}_{y}$. Binary values are obtained with the CLIP model [44] by comparing the similarity of each image to the embedding of an attribute and to its negative counterpart.
@@
The Caltech-UCSD Birds (CUB) dataset contains 11,788 images of 200 bird species, annotated with part locations, bounding boxes, and 312 binary attributes. Following the approach of Koh et al. [9], we retain only the attributes that are present in at least 10 species (based on majority voting), resulting in a filtered set of 112 attributes. These attributes are grouped into 27 semantic clusters, where each group is defined by a common prefix in the attribute names. In our setup, the task variable $\bm{y}$ is to classify the bird species. The concept set $\bm{c}$ consists of the attributes belonging to 12 randomly selected groups (per run), while the nuisance set $\bm{n}y$ includes the attributes from the remaining 20 groups. Since most attributes exhibit some correlation with the classification task, we set $\bm{n}_{\bar{y}}$ to the empty set.